### v1

# A Fast Kernel-based Conditional Independence test with Application to Causal Discovery

Abstract
Kernel-based conditional independence (KCI) testing is a powerful nonparametric method commonly employed in causal discovery tasks. Despite its flexibility and statistical reliability, cubic computational complexity limits its application to large datasets. To address this computational bottleneck, we propose FastKCI, a scalable and parallelizable kernel-based conditional independence test that utilizes a mixture-of-experts approach inspired by embarrassingly parallel inference techniques for Gaussian processes. By partitioning the dataset based on a Gaussian mixture model over the conditioning variables, FastKCI conducts local KCI tests in parallel, aggregating the results using an importance-weighted sampling scheme. Experiments on synthetic datasets and benchmarks on real-world production data validate that FastKCI maintains the statistical power of the original KCI test while achieving substantial computational speedups. FastKCI thus represents a practical and efficient solution for conditional independence testing in causal inference on large-scale data.
1 Introduction
Conditional independence (CI) testing is a fundamental operation in causal discovery and structure learning. Widely used algorithms such as the PC algorithm [25] and Fast Causal Inference [24] rely on CI tests to recover the causal skeleton of a graph from observational data. The core statistical question is whether two variables $X$ and $Y$ are independent given a conditioning set $Z$, that is, whether $X\perp\!\!\!\!\perp Y\mid Z$. Despite being frequently adapted to fields like neuroscience [23], climate research [7] or economics [2], CI testing remains the computational bottleneck in constraint-based causal discovery, especially as sample size increases [1, 17, 22].
Standard CI tests suit different types of data and assumptions: Traditional tests like the Fisher-$Z$ test [8] assume linear Gaussian data, while discrete tests such as $\chi^{2}$ require categorical variables. More recent approaches avoid strong distributional assumptions by leveraging nonparametric techniques such as kernel methods. Among these, the Kernel-based Conditional Independence test (KCI) [32] has become a standard choice due to its flexibility and empirical power. KCI is based on Hilbert space embeddings and computes dependence via kernel covariance operators, making it applicable to arbitrary continuous distributions.
KCI requires operations on $n\times n$ Gram matrices and matrix inversions, resulting in $\mathcal{O}(n^{3})$ runtime per test, which makes it infeasible for large-scale applications. Recent work has sought to mitigate this cost via sample splitting [19], random Fourier features [28], neural network approximations [6], and randomization-based tests [21]. However, these approximations can degrade statistical power and require additional tuning.
Our goal is to accelerate the KCI test without sacrificing its statistical rigor and nonparametric flexibility. To this end, we propose FastKCI, a novel variant that leverages ideas from embarrassingly parallel inference in Gaussian processes [34]. FastKCI partitions data based on a generative model in the conditioning set $Z$, performs KCI tests in parallel on each partition, and aggregates the test statistics using an importance weighting scheme. This blockwise strategy enables significant computational speedups, especially in multi-core or distributed environments.
Our contributions are therefore a scalable and parallelizable conditional independence test, FastKCI, that significantly accelerates KCI by using a novel partition-based strategy combined with importance weighting. We experimentally demonstrate that FastKCI retains the statistical performance of KCI while achieving runtime improvements across synthetic and real-world datasets.
2 Related Work
Due to the growing importance of causal inference, considerable research has been devoted to discovering efficient methods to identify causal structures from observational data [29]. Traditional causal discovery methods are often categorized as constraint-based or score-based. Constraint-based methods, such as the PC algorithm [25] or FCI [26], rely on CI tests to identify the underlying causal structure. Score-based methods, such as Greedy Equivalence Search [3], evaluate causal structures based on scoring criteria. Both approaches face computational challenges: constraint-based methods due to intensive CI testing, and score-based methods due to an exponentially large search space.
Existing work has improved the efficiency of the PC algorithm by reducing unnecessary CI tests [27], optimizing the overall search procedure itself—through order-independent execution [4], skipping costly orientation steps [5], sparsity-aware pruning [15], divide-and-conquer partitioning [13], and parallelism of the CI tests [17, 30, 12]. These methods, however, do not address the cubic-time bottleneck of kernel-based CI tests.
Attempts in the literature to address this issue take several forms. The original KCI paper already derived an analytic $\Gamma$-approximation of the null distribution and proposed simple median-heuristic bandwidth choices to avoid costly resampling [32]. Strobl et al. [28] accelerate KCIT by replacing the full kernel matrices with an $m$-dimensional random Fourier-feature approximation. Doran et al. [6] re-express conditional independence as a single kernel two-sample problem by restricting permutations of $(X,Y)$, thereby changing the test statistic while lowering runtime. Zhang et al. [31] eliminate kernel eigen-decompositions altogether by regressing $X$ and $Y$ on $Z$ and measuring residual similarity with a lightweight kernel. Additional ideas include calibrating test statistics with locality-based permutations in the conditioning set [16], evaluating analytic kernel embeddings at a finite set of landmark points [20], and controlling small-sample bias via data splitting [19]. However, these approaches can compromise statistical power, particularly under complex nonlinear dependencies and in high-dimensional conditioning sets.
Our method preserves the KCI statistic by evaluating it on Gaussian-mixture strata and aggregating with importance weights. We therefore leverage techniques from Gaussian Process regression, which often faces the identical problem of poor scalability due to cubic complexity [^1] In fact, as KCI solves GP regression problems to find the RKHS bandwidth, it is a sub-problem of CI. [34]. A common solution is to assume the underlying distribution of the covariates $Z=\{z_{1},\dots,z_{j}\}$ to be a mixture-of-experts (MoE) [14]. Local approaches, as in Gramacy and Lee [10], then aggregate over multiple partitions by MCMC. Zhang and Williamson [34] propose an importance sampling approach to MoE that efficiently aggregates over multiple partitions. To the best of our knowledge, our approach—using importance-sampled partitions of the data, performing parallel kernel tests, and aggregating via importance weighting—represents the first attempt explicitly bridging embarrassingly parallel inference with scalable kernel-based CI testing.
3 Background
The KCI builds on a notion of conditional independence introduced by Fukumizu et al. [9]. Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space and $X\in\mathcal{X}$, $Y\in\mathcal{Y}$, $Z\in\mathcal{Z}$ random variables. For each domain we fix a measurable, bounded and characteristic kernel, e.g. the Gaussian RBF, denoted $k_{X},k_{Y},k_{Z}$, with corresponding reproducing-kernel Hilbert spaces (RKHS) $\mathcal{H}_{X},\mathcal{H}_{Y},\mathcal{H}_{Z}$. Feature maps are denoted $\varphi_{X}(x)=k_{X}(x,\cdot)$ etc. Expectations are shorthand $\mathbb{E}[\cdot]$, tensor products $\otimes$, and centered features $\tilde{\varphi}_{X}\!:=\!\varphi_{X}-\mu_{X}$ where $\mu_{X}=\mathbb{E}[\varphi_{X}(X)]$.
3.1 Covariance and conditional covariance operators
The cross-covariance operator $\Sigma_{XY}\colon\mathcal{H}_{X}\!\to\!\mathcal{H}_{Y}$ is the bounded linear map satisfying
$\langle g,\Sigma_{XY}f\rangle=\mathbb{E}\!\bigl{[}\langle\tilde{\varphi}_{X}(X),f\rangle_{\!X}\;\langle\tilde{\varphi}_{Y}(Y),g\rangle_{\!Y}\bigr{]}\quad\forall\,f\!\in\!\mathcal{H}_{X},\,g\!\in\!\mathcal{H}_{Y}.$
An analogous definition yields $\Sigma_{XZ}$, $\Sigma_{ZZ}$. Provided $\Sigma_{ZZ}$ is injective [^2] With a characteristic kernel this holds on the closure of its range., conditional covariance is
$\Sigma_{XY\mid Z}:=\Sigma_{XY}-\Sigma_{XZ}\,\Sigma_{ZZ}^{-1}\Sigma_{ZY}.$
Proposition 1 (Fukumizu et al., 2007).
With characteristic kernels, $X\perp\!\!\!\!\perp Y\mid Z\Longleftrightarrow\Sigma_{XY\mid Z}=0.$
Hence testing conditional independence reduces to checking whether this operator is null.
3.2 Finite-sample KCI statistic
Given $n$ observations $\{(x_{i},y_{i},z_{i})\}_{i=1}^{n}$, assemble Gram matrices $K_{X},K_{Y},K_{Z}\!\in\!\mathbb{R}^{n\times n}$ with $(K_{X})_{ij}=k_{X}(x_{i},x_{j})$, $K_{Y},K_{Z}$ analogously. Let $H:=I_{n}-\tfrac{1}{n}\mathbf{1}\mathbf{1}^{\!\top}$ and define the projection onto $Z$–residuals
$R_{Z}:=I_{n}-K_{Z}\bigl{(}K_{Z}+\lambda I_{n}\bigr{)}^{-1},\quad\lambda>0.$ (1)
Intuitively, $R_{Z}$ acts like a regression operator that projects any function of the samples onto the subspace of functions depending on $Z$. Thus, $I-R_{Z}$ removes components explained by $Z$. Residualized, centered kernels are $\tilde{K}_{X}=R_{Z}HK_{X}HR_{Z}$ and $\tilde{K}_{Y}$ analogously.
The KCI test statistic of Zhang et al. [32] is then given by the Hilbert–Schmidt norm of the cross-covariance between residuals:
$T_{\text{KCI}}:=\frac{1}{n}\operatorname{Tr}(\tilde{K}_{X}\tilde{K}_{Y}).$ (2)
Zhang et al. [32] show that under $H_{0}$ this statistic converges in distribution to a weighted sum of $\chi^{2}$ variables. In practice one uses a finite-sample null distribution to assess significance. We follow Proposition 5 in [32], using a spectral approach to simulate the null: We compute the eigenvalues ${\lambda_{m}}$ of a normalized covariance operator associated with $\tilde{K}_{X}$ and $\tilde{K}_{Y}$, then generate null samples and set
$T_{\text{null}}^{(b)}=\sum_{m=1}^{n}\lambda_{m}\,\chi_{1,m,b}^{2},\qquad b=1,\dots,B,$ (3)
with i.i.d. $\chi_{1}^{2}$ variates $\chi_{1,m,b}^{2}$. This approximate distribution of $T_{\text{KCI}}$ under $H_{0}$ obtains a $p$-value. The full KCI procedure is given in 2.
3.3 Computational Complexity of the KCI
A key drawback of KCI is its heavy computation for large $n$. Constructing and manipulating $n\times n$ Gram matrices is $\mathcal{O}(n^{2})$ in memory. More critically, forming $\tilde{K}_{X}$ and $\tilde{K}_{Y}$ requires solving a linear system or eigen-decomposition on $K_{Z}\in\mathbb{R}^{n\times n}$. The inversion $(K_{Z}+\varepsilon I)^{-1}$ costs $\mathcal{O}(n^{3})$ time in general. Generating the null distribution via eigenvalues also incurs an $\mathcal{O}(n^{3})$ decomposition of the $n\times n$ matrix $U=(I-R_{Z})K_{X}(I-R_{Z})$ and related matrices. Thus, the overall complexity of KCI scales cubically in the sample size $n$. This cubic bottleneck severely limits the test’s applicability to large datasets, particularly when it must be repeated many times as in constraint-based causal discovery.
4 FastKCI: A Scalable and Parallel Kernel-based Conditional Independence Test
To overcome the $\mathcal{O}(n^{3})$ bottleneck, we propose FastKCI, which leverages a mixture-of-experts model and importance sampling. The core idea is to break the full kernel computation into $V$ smaller pieces (experts) corresponding to data partitions, compute local CI statistics on each piece, and then recombine them to recover the global statistic. By doing so, FastKCI achieves significant speed-ups – roughly on the order of $1/V^{2}$ of the cost of KCI – while provably maintaining the test’s correctness under mild assumptions.
We assume that the distribution of the conditioning variable $Z$ can be approximated by a mixture of $V$ Gaussian components.
Assumption 1.
Let $U$ be a latent cluster assignment variable taking values in $\{1,\dots,V\}$. For each $i$, we posit a model:
$U_{i}\sim\text{Categorical}(\pi_{1},\dots,\pi_{V}),\quad\pi\sim\text{Dirichlet}(\alpha),$ (4)
$z_{i}\mid(U_{i}=v)\sim\mathcal{N}(\mu^{Z}_{v},\Sigma^{Z}_{v}),\quad(\mu^{Z}_{v},\Sigma^{Z}_{v})\sim\text{Normal-InvWishart}(\mu_{0},\lambda_{0},\Psi,\nu).$ (5)
Thus, ${z_{i}}$ are i.i.d. draws from a $V$-component Gaussian mixture with unknown means $\mu_{v}^{Z}$ and covariances $\Sigma_{v}^{Z}$. We place a weak Normal-Inverse-Wishart prior on these parameters to allow uncertainty. This mixture-of-experts prior on $Z$ guides partitioning of the dataset. Intuitively, if $Z$ has a multimodal or complex distribution, this model lets us divide the data into $V$ clusters $\mathcal{C}_{1},\dots,\mathcal{C}_{V}$ (with $\mathcal{C}_{v}=\{i\mid U_{i}=v\}$) such that within each $\mathcal{C}_{v}$, $Z$ is roughly Gaussian. We do not assume anything restrictive about $X$ and $Y$ globally. By conditioning on $\mathcal{C}_{v}$, the relationship between $X$ and $Y$ can be analyzed locally. In particular, conditioned on a given partition $U_{1:n}=(U_{1},\dots,U_{n})$, the kernel matrices $K_{X},K_{Y},K_{Z}$ acquire an approximate block structure: after permuting indices by cluster, each matrix breaks into $V$ blocks (sub-matrices) corresponding to points in the same cluster, with negligible entries for cross-cluster pairs, especially if clusters are well-separated in $Z$. Each cluster $v$ defines a local sub-problem of size $n_{v}=|\mathcal{C}_{v}|$, and within that cluster we can perform a CI test on the restricted data $\{(x_{i},y_{i},z_{i}):i\in\mathcal{C}_{v}\}$. This yields a local test statistic $T_{v}$ and local null distribution for cluster $v$. By appropriately combining these local results, we can recover a valid global test statistic without ever computing the full $n\times n$ kernel on all data at once. In the next paragraphs, we explain the procedure in detail.
Figure 1: Motivation of the partitioning scheme in the data. The full covariance kernel estimation (left) is inefficient, while creating a partition (middle) may neglect some covariance structure. We propose to use multiple partitions (right) in parallel across $J$ parallel cores. We combine them using importance sampling. The figure is inspired by Zhang [33].
Partition Sampling.
As visible in Figure 1, for $j=1,\dots,J$ we independently draw a cluster assignment $U^{(j)}_{1:n}=(U_{1}^{(j)},\dots,U_{n}^{(j)})$ according to the MoE model on $Z$. The mixture prior is fit to the empirical $Z$ data. For example, one iteration $j$ samples mixture parameters $(\mu_{1}^{Z},\dots,\mu_{V}^{Z},\Sigma_{1}^{Z},\dots,\Sigma_{V}^{Z})\sim P(\mu,\Sigma)$ from the NIW prior and mixture weights $\pi$, then assigns each point $i$ to a cluster $U_{i}^{(j)}=v$ with probability
$P\left(U_{i}^{(j)}=v\mid z_{i}\right)\propto\pi_{v}\mathcal{N}\left(z_{i}\mid\mu_{v}^{Z},\Sigma_{v}^{Z}\right).$
Each such draw yields a partition $\{\mathcal{C}_{1}^{(j)},\dots,\mathcal{C}_{V}^{(j)}\}$ of $\{1,\dots,n\}$.
Local RKHS Embedding, Test Statistic and Null Distribution.
For each partition $j$ and for each cluster $v\in\{1,\dots,V\}$, let $\mathcal{C}_{v}^{(j)}=\{i\mid U_{i}^{(j)}=v\}$ be indices in the partition of size $n_{v}^{(j)}=|\mathcal{C}_{v}^{(j)}|$. We form $n_{v}^{(j)}\times n_{v}^{(j)}$ Gram matrices $K_{X}^{(j,v)}$, $K_{Y}^{(j,v)}$, and $K_{Z}^{(j,v)}$. $\tilde{K}_{X}^{(j,v)}$ and $\tilde{K}_{Y}^{(j,v)}$ are calculated analogously to Equation 1 with local regression operators. The local test statistic follows as
$T_{v}^{(j)}=\frac{1}{n_{v}^{(j)}}\mathrm{Tr}\left(\tilde{K}_{X}^{(j,v)}\tilde{K}_{Y}^{(j,v)}\right).$
We generate a set of $B$ null samples $\{T_{v,\text{null}}^{(j,b)}\}$ for each cluster by applying the spectral method in Equation 3 block-wise. Finally, we record a log-likelihood score for each cluster: let $\mathcal{L}\left(X^{(j,v)}\right)=\log P\left(X\left(\mathcal{C}_{v}^{(j)}\right)\mid Z\left(\mathcal{C}_{v}^{(j)}\right)\right)$ and $\mathcal{L}\left(Y^{(j,v)}\right)$ analogously be the log marginal likelihoods of the $X$ and $Y$ data in cluster $v$ given $Z$. We use this measure to calculate a likelihood $\ell^{(j,v)}\propto P(X^{(j)},Y^{(j)}\mid U^{(j)})$.
Aggregation over $V$.
We aggregate the partition-wide test statistic as the sum of the cluster statistics.
$T^{(j)}=\sum_{v}T_{v}^{(j)}$
This recovers the full trace of the product of block-wise diagonal $\tilde{K}_{x}^{(j)}\tilde{K}_{y}^{(j)}$. Since under $H_{0}$ the cluster test statistics are approximately independent (different clusters involve disjoint data) and each follows a weighted $\chi^{2}$ distribution, their sum $T_{\text{null}}^{(j,b)}$ is a valid sample from the null distribution for the whole partition $j$. Thus, we also aggregate each null sample into the sum.
$T_{\mathrm{null}}^{(j,b)}=\sum_{v}T_{v,\mathrm{null}}^{(j,b)}$
Aggregation over $J$.
We apply a softmax to the log-likelihoods to obtain per-partition importance weights
$w_{j}=\frac{\exp\left(\sum_{v}\ell^{(j,v)}\right)}{\sum_{j}\exp\left(\sum_{v}\ell^{(j,v)}\right)}$
The final test statistic is a weighted average across all $J$ partitions:
$T_{\text{FastKCI}}=\sum_{j=1}^{J}w_{j}T^{(j)},\quad\text{where }w_{j}\propto P(X^{(j)},Y^{(j)}\mid U^{(j)})$
and the combined null distribution is taken as the mixture of all partition null samples with the same weights. In practice, we merge the $J$ sets of null samples $\{T_{\text{null}}^{(j,b)}\}$ into one weighted empirical distribution. Specifically, we compute the weighted empirical cumulative density $F_{T,\text{null}}(t)=\sum_{j=1}^{J}w_{j}\Big{(}\frac{1}{B}\sum_{b=1}^{B}\mathbf{1}\{T_{\text{null}}^{(j,b)}\leq t\}\Big{)}$, which is a mixture of the $J$ null distributions.
4.1 Theoretical Insight
Under the null hypothesis $H_{0}:X\perp\!\!\!\!\perp Y\mid Z$, each block statistic $T^{(j)}_{v}$ is, by exactly the same argument as in [32], asymptotically a weighted sum of independent $\chi^{2}_{1}$ variables,
[…]
with non-negative weights $\lambda_{j,v,m}$ determined by the eigenvalues of the blockwise covariance operators.
Different clusters use disjoint subsets of the data, hence their statistics are (asymptotically) independent; the sum $T^{(j)}=\sum_{v}T^{(j)}_{v}$ is therefore a new weighted $\chi^{2}$ mixture whose weights are a union of $\lambda_{j,v,m}$. The convex combination of these weighted $\chi^{2}$ mixtures is itself a weighted $\chi^{2}$ mixture with the same weights $w_{j}$ :
[…]
Hence, our method inherits exactly the same null-law template as classical KCI: a positive, finite linear combination of $\chi^{2}_{1}$ variables. [^3] The formal requirements are the standard KCI assumptions (characteristic kernels, boundedness) plus every cluster size $|\mathcal{C}^{(j)}_{v}|\!\to\!\infty$ as $n\to\infty$. Consequently, under Assumption 1 and with $J\rightarrow\infty$, the test exhibits the same appealing statistical properties as the conventional KCI.
4.2 Computational Complexity
Assuming balanced partitions, each block contains roughly $n/V$ samples. The complexity of KCI per block is then $\mathcal{O}((n/V)^{3})$, and the total complexity becomes $\mathcal{O}(Jn^{3}/V^{2})$. Since each of the $J$ runs is fully parallelizable, the wall-clock cost is significantly reduced compared to the original $\mathcal{O}(n^{3})$ cost.
The complete procedure is summarized in Algorithm 1.
Algorithm 1 FastKCI: Fast and Parallel Kernel-based CI Test
0:  Dataset $\{(x_{i},y_{i},z_{i})\}_{i=1}^{n}$, number of partitions $V$, number of samples $J$.
1:  for $j=1$ to $J$ in parallel do
2:     Sample $V$-component partition $U^{(j)}$ from $p(U\mid Z)$.
3:     for each component $v=1$ to $V$ do
4:        Compute residualized kernels $K_{X|Z}^{(v)}$, $K_{Y|Z}^{(v)}$.
5:        Compute test statistic $T_{v}^{(j)}=\frac{1}{n_{v}}\mathrm{Tr}(K_{X|Z}^{(v)}K_{Y|Z}^{(v)})$.
6:     end for
7:     Aggregate: $T^{(j)}=\sum_{v}T_{v}^{(j)}$.
8:     Compute importance weight $w_{j}\propto P(X^{(j)},Y^{(j)}\mid U^{(j)})$.
9:  end for
10:  Normalize weights: $w_{j}\leftarrow w_{j}/\sum_{j}w_{j}$.
11:  Compute final test statistic and null distribution by weighted averaging.
12:  return  $p$-value for $H_{0}:X\perp\!\!\!\!\perp Y\mid Z$.
5 Experiments
To empirically prove the main result in Algorithm 1, we extensively study the performance of FastKCI and compare it to the KCI implementation provided in the causal-learn package [35]. We consider different scenarios, focusing on coverage, power and causal discovery. [^4] Our full code can be found and used within the causal-learn package in python. Please see https://github.com/py-why/causal-learn/blob/main/causallearn/utils/FastKCI/FastKCI.py.
5.1 Coverage Comparison
We run experiments inspired by Zhang et al. [32], namely, we generate $n=1200$ samples of random variables $X$, $Y$ and $Z$, with $X$ and $Y$ being drawn independently conditioned on $Z$. In our scenario, we examine the type I error with growing confounding set size $D=\{1,\dots,5\}$, with all variables effecting both $X$ and $Y$ (comparable to “Case II” in Zhang et al. [32]). For the $Z_{i}$, we consider multiple ground-truth distributions, as a mixture of $V_{\text{true}}=\{1,3,10\}$ Gaussians. $X$ and $Y$ are generated as $g\left(\sum_{i}f_{i}(Z_{i})+\varepsilon\right)$ where $f$ and $g$ are random mixtures of linear, cubic and tanh functions and $\varepsilon$ is independent across $X$ and $Y$.
Figure 2: Simulated coverage of the FastKCI and the KCI at 1% and 5% levels.
5.2 Power Comparison
We repeat the experiments from above but make $X$ and $Y$ conditionally dependent by adding a small identical noise component $\nu\sim\mathcal{N}(0,\sigma^{2}_{\textrm{vio}})$ to both of them and assess the type II error. We compare KCI and FastKCI in different configurations with a growing violation of $H_{0}$ (i.e., $\sigma^{2}_{\textrm{vio}}$ is increasing). Figure 3 displays that both approaches have similar performance at a sample size of $n=1200$.
Figure 3: Power of FastKCI in different configurations compared to KCI. The violation of the null-hypothesis is increasing on the x-axis.
$V_{\text{true}}$ Algorithmn Power ($\alpha=5\%$) Power ($\alpha=1\%$)
1 FastKCI(V=10, J=128) 0.67 0.43
FastKCI(V=10, J=16) 0.665 0.405
FastKCI(V=3, J=128) 0.86 0.745
FastKCI(V=3, J=16) 0.87 0.665
KCI 0.91 0.77
3 FastKCI(V=10, J=128) 0.725 0.545
FastKCI(V=10, J=16) 0.79 0.62
FastKCI(V=3, J=128) 0.905 0.74
FastKCI(V=3, J=16) 0.89 0.735
KCI 0.86 0.72
10 FastKCI(V=10, J=128) 0.73 0.55
FastKCI(V=10, J=16) 0.755 0.54
FastKCI(V=3, J=128) 0.87 0.645
FastKCI(V=3, J=16) 0.82 0.665
KCI 0.855 0.605
Table 1: Power of KCI and FastKCI under violation of $H_{0}$. $X$ and $Y$ are confounded by $Z$, but there is also a direct edge between them.
For another comparison, we consider the setting from Section 5.1, but now $X$ directly causes $Y$. We calibrate the setting to a small violation, under which we observe a non-zero type II error. [^5] The magnitude of the violation is approximately $1/3$ of the signal. Table 1 shows the power under different numbers of Gaussians $V_{\text{true}}$ in the DGP.
5.3 Causal Discovery
We compare the performance of the PC algorithm using FastKCI with KCI in causal discovery tasks. For this, we consider two different settings, setting A is derived from Zhang et al. [32]. We sample $6$ random variables $\{X_{1},\dots,X_{6}\}$. For $j>i$ we sample edges with probability $0.3$. Based on the resulting DAG, we sample descendants from a Gaussian Process with mean function $\sum_{i\in\text{Pa}(X_{j})}\nu_{i}\cdot X_{i}$ (with $\nu_{i}\sim\mathcal{U}[-2,2]$) and a covariance kernel consisting of a Gaussian kernel plus a noise kernel. Setting B is derived from Liu et al. [18] and similarly consists of $6$ random variables. For $j>i$, we sample edges with probability $0.5$. The link function $f_{i}$ is randomly chosen between being linear and non-linear. For linear components, the edge weights are drawn from $\mathcal{U}[-1.5,-0.5],[0.5,1.5]$, while non-linear components follow multiple functions (sin, cos, tanh, sigmoid,polynomial). Added noise is simulated from $\mathcal{N}(0,\sigma^{2}_{i})$ with $\sigma\in\{0.2,0.5\}$. As shown in figure 4, both methods exhibit similar performance in precision, recall and F1 score of the discovered edges.
Figure 4: Precision, recall and F1-Score for KCI and FastKCI in causal discovery with growing sample size in setting A (left) and setting B (right).
5.4 Scalability
As Figure 5 highlights, FastKCI shows excellent scalability, particularly when using a large number of blocks $V$. To further investigate the scalability of FastKCI, we scale up the sample size in the experiments and report precision, recall, f1 and computation time. [^6] To reduce complexity, in these experiments we approximate the kernel bandwidth instead of determinating it exactly with GP. See [32] for detail. The results in Table 2 show that we achieve good results in feasible time even for sample sizes where the traditional KCI is fails due to memory and CPU constraints.
DGP n Method Precision Recall F1 Time [s]
Gaussian Process 2000 KCI 0.9833 1.0000 0.9910 467.48
FastKCI(V=3) 0.9526 1.0000 0.9740 250.59
FastKCI(V=10) 0.8980 0.9500 0.9213 129.40
10000 KCI 0.9130 1.0000 0.9496 22240
FastKCI(V=3) 0.9130 1.0000 0.9496 9136.5
FastKCI(V=10) 0.9167 1.0000 0.9524 1505.2
Nonlinear Process 2000 KCI 0.9704 0.9299 0.9470 1098.9
FastKCI(V=3) 0.9783 0.9251 0.9484 587.72
FastKCI(V=10) 0.9641 0.8861 0.9196 219.73
10000 KCI 1.0000 1.0000 1.0000 99360
FastKCI(V=3) 1.0000 0.9711 0.9847 51709
FastKCI(V=10) 1.0000 0.9841 0.9916 12152
20000 FastKCI(V=10) 1.0000 1.0000 1.0000 68086
FastKCI(V=50) 0.9500 0.9722 0.9575 1909.7
50000 FastKCI(V=50) 0.9667 0.9818 0.9739 74095
FastKCI(V=100) 0.9818 0.9533 0.9664 7959.4
100000 FastKCI(V=100) 1.0000 1.0000 1.0000 86342
FastKCI(V=200) 0.8611 1.0000 0.9251 12253
Table 2: Precision, recall, F1 and computational time of the KCI and FastKCI on very large samples.
5.5 Computation Time
We showcased the empirical performance of FastKCI for both pure CI tasks as well as causal discovery with PC. To shed light on the computational speed-up, we report the computation time in Figure 5), which – depending on the choice of the main tuning parameters $V$ and $J$ – is significantly faster than for the KCI.
Figure 5: Computation time with increasing sample size for (a) conditional independence testing and (b) causal discovery with the PC algorithm.
6 Application to Production Data
We apply our proposed method to a semi-synthetic dataset for causal discovery, contained in causal-assembly [11]. The ground-truth consists of $98$ production stations, each dedicated to specific automated manufacturing processes where individual components are progressively assembled. The processes involved, such as press-in and staking, are mechanically complex and non-linear. The resulting data provides a real-world example on which causal discovery can enhance the understanding of causes for production results and yields a good benchmark for our proposed methodology.
We compare KCI with the FastKCI variant, setting $V=10$. Table 3 depicts the full assembly line with $98$ nodes, while Table 4 shows results only for one of the production stations with $16$ nodes. Please refer to Göbler et al. [11] for an overview of results with alternative CD algorithms such as PC with fisher-Z, lingam and others. In terms of precision and recall, KCI performs slightly better than FastKCI, but both outperform all tested methods in the original benchmark paper. In terms of computational speed, FastKCI, especially in the station setting with a smaller number of nodes, has a significant edge over the standard method.
Type Mean Precision Mean Recall Mean F1 Mean Computation Time [s]
PC (Fast KCI) 0.6263 0.0951 0.1651 21107
PC (KCI) 0.5894 0.1325 0.2163 24659
snr 0.3501 0.1311 0.0926
grandag 0.3193 0.0109 0.0045
lingam 0.3281 0.1092 0.0721
PC (Fisher-Z) 0.4121 0.1170 0.0968
notears 0.5209 0.0978 0.1019
das 0.2971 0.0784 0.0474
Table 3: Mean result of $10$ repetitions on the Causal Assembly benchmark with $n=1000$.
Type Mean Precision Mean Recall Mean F1 Mean Computation Time [s]
Fast KCI (V=10) 0.7815 0.8167 0.7982 371.5
KCI 0.7751 0.8847 0.8249 1230.1
Table 4: Mean result of $100$ repetitions on the Causal Assembly benchmark, Station 3, with $n=2000$.
7 Conclusion and Limitations
Our paper introduced FastKCI, which turns the cubic-time KCI into a more scalable and embarrassingly parallel procedure. The clustering approach in the conditioning set $Z$ via a MoE assumption is unbiased under mild conditions, while the per test cost falls roughly by a factor $1/V^{2}$. Our experiments show that FastKCI preserves KCI’s Type-I and Type-II error across diverse settings and cuts wall-clock time for both CI and causal discovery dramatically as sample size grows. Thus, our method is a promising approach for scaling up CI testing.
The main limitation is reliance on the Gaussian-mixture assumption for partitioning, although is is a common assumption for multimodal settings, misleading clusters can hurt power, particularly under small $J$. Future work should include exploring alternative sample partitioning and importance weighting schemes in order to further refine efficiency of FastKCI.
Appendix A Appendix
A.1 Additional Results
Here we provide additional results for the other stations of the causal assembly benchmark.
Type Mean Precision Mean Recall Mean F1 Mean Computation Time [s]
Fast KCI (V=10) 1.000 0.4314 0.5938 25.369
KCI 0.9891 0.6814 0.7994 235.12
Table 5: Mean result of $100$ repetitions on the causal assembly benchmark, Station 1, with $n=2000$.
Type Mean Precision Mean Recall Mean F1 Mean Computation Time [s]
Fast KCI (V=10) 0.6239 0.3240 0.4262 4427.1
KCI 0.5963 0.3980 0.4772 25279
Table 6: Mean result of $100$ repetitions on the causal assembly benchmark, Station 2, with $n=2000$.
Type Mean Precision Mean Recall Mean F1 Mean Computation Time [s]
Fast KCI (V=10) 0.5236 0.3499 0.4190 3282.8
KCI 0.5536 0.4665 0.5057 21755
Table 7: Mean result of $100$ repetitions on the causal assembly benchmark, Station 4, with $n=2000$.
Type Mean Precision Mean Recall Mean F1 Mean Computation Time [s]
Fast KCI (V=10) 0.4619 0.5350 0.4953 543.38
KCI 0.4822 0.6130 0.5390 2550.6
Table 8: Mean result of $100$ repetitions on the causal assembly benchmark, Station 5, with $n=2000$.
A.2 Kernel-based Conditional Independence Test (KCI)
Here we provide a clearly structured algorithm box summarizing the original Kernel-based Conditional Independence (KCI) test introduced by Zhang et al. [32] for clarity and comparison with the fast procedure.
Algorithm 2 Kernel-based Conditional Independence Test (KCI) by Zhang et al. [32]
0:  Datasets $\{x_{i},y_{i},z_{i}\}_{i=1}^{n}$ where $x_{i}\in\mathbb{R}^{d_{x}}$, $y_{i}\in\mathbb{R}^{d_{y}}$, $z_{i}\in\mathbb{R}^{d_{z}}$; kernel functions $k_{x}$, $k_{y}$, and $k_{z}$ for $X$, $Y$, and $Z$ respectively; regularization parameter $\lambda$.
0:  Test decision for $H_{0}:X\perp Y\mid Z$
1:  Aggregate $\ddot{X}=(X,Z)$. Compute the kernel matrices $K_{\ddot{X}}$, $K_{Y}$, and $K_{Z}$ for the datasets using the respective kernel functions:
$(K_{\ddot{X}})_{ij}=k_{x}({\ddot{x}}_{i},\ddot{x}_{j}),\quad(K_{Y})_{ij}=k_{y}(y_{i},y_{j}),\quad(K_{Z})_{ij}=k_{z}(z_{i},z_{j})$
2:  Center the kernel matrices $K_{\ddot{X}}$, $K_{Y}$ and $K_{Z}$:
$\tilde{K}_{\ddot{X}}=HK_{\ddot{X}}H,\quad\tilde{K}_{Y}=HK_{Y}H\quad\tilde{K}_{Z}=HK_{Z}H$
where $H=I-\frac{1}{n}\mathbf{1}\mathbf{1}^{T}$ and $\mathbf{1}$ is a column vector of ones.
3:  Calculate the projection matrix from a kernel ridge regression on $Z$:
$R_{Z}=I-\tilde{K}_{Z}(\tilde{K}_{Z}+\lambda I)^{-1}=\lambda(\tilde{K}_{Z}+\lambda I)^{-1}$
4:  Compute the residual kernels of $\ddot{X}$ and $Y$ after conditioning on $Z$:
$\tilde{K}_{\ddot{X}\mid Z}=R_{Z}\tilde{K}_{X}R_{Z},\quad\tilde{K}_{Y\mid Z}=R_{Z}\tilde{K}_{Y}R_{Z}$
5:  Compute the test statistic:
$T_{CI}\triangleq\frac{1}{n}\text{Tr}\left(\tilde{K}_{\ddot{X}\mid Z}\tilde{K}_{Y\mid Z}\right)$
6:  Bootstrap samples from the asymptotic distribution of the test statistic $\check{t}_{CI}$ [^7] For simplicity, we refer for details to Zhang et al. [32]. Sampling from the null distribution involves multiple eigendecompositions.
7:  Compute the p-value:
$p=\frac{1}{B}\sum_{b=1}^{B}\mathbb{I}(\check{t}_{CI}^{b}\geq T_{CI})$
8:  return  p-value for $H_{0}$
77footnotetext: For simplicity, we refer for details to Zhang et al. [32]. Sampling from the null distribution involves multiple eigenvalue and vector computations.
Appendix B Computing Resources.
[…]